OpenStudy (anonymous):

How do I find ln e^3

OpenStudy (anonymous):

calculator i think

OpenStudy (anonymous):

e^3 is around 20.09

OpenStudy (anonymous):

also if you set it into a logarithm: ln(x) = 3

OpenStudy (helder_edwin):

\[\large \ln e^3=3\ln e=3 \]